OpenStudy (anonymous):

if the bullet is fired for maximum horizontal range of 2560m with a speed of 160m/s find the maximum the bullet can reach it (g=10m/s^2)

OpenStudy (anonymous):

maximum what?

OpenStudy (experimentx):

You have maximum range when angle of projection is equal to 45. horizontal velocity = \( u \cos \theta \) Distance travelled = velocity x time = \( t \times u \cos \theta \) = 2560m vertical velocity = \( u \times \sin\theta\) since you know time ... find the range

OpenStudy (anonymous):

for maximum range, it should be fired at an angle of 45 degrees \[R = u ^{2}\sin 2\theta /g\]

OpenStudy (anonymous):

we have to find the maximum height>>>

OpenStudy (ujjwal):

what do you mean by " find the maximum the bullet can reach it " .. I didn't get the question.. Are you sure you haven't made typing errors?

OpenStudy (anonymous):

sorry for the mistake

OpenStudy (anonymous):

it is maximum height...

OpenStudy (ujjwal):

I see.. max height .. wait a moment..

OpenStudy (anonymous):

ok

OpenStudy (ujjwal):

Maximum horizontal range is for angle of projection 45 degrees.. Now in formula of R put theta = 45 and find u .. Now replace value of u and theta=45 in formula of H.. and you will get it..

OpenStudy (anonymous):

we are given the range. so find the angle thete. it comes out to be 45 degrees. now substitute these values in the max. height equation. \[H = u ^{2} \sin ^{2}\theta/2g\]

OpenStudy (anonymous):

@ujjwal i think we are given the value of initial velocity (u)

OpenStudy (anonymous):

yes u=160m/s

OpenStudy (anonymous):

go ahead and calculate the max. height

OpenStudy (anonymous):

i am getting it as 320 but the answer is 640m

OpenStudy (ujjwal):

It is a lot easier than.. you have more info than needed.. go ahead put theta = 45 .. that's all.. and i recommend you to go through the theory part of your text book before you dive into numericals...

OpenStudy (anonymous):

a calculation mistake

OpenStudy (anonymous):

yea i got theta =30

OpenStudy (anonymous):

how u got theta =45 can u show me

OpenStudy (anonymous):

\[2560 = (16)^{2}\times \sin2\theta /10\]

OpenStudy (anonymous):

i'm sorry its (160)^2

OpenStudy (anonymous):

sin 2 theta =1 then what to do???

OpenStudy (ujjwal):

And it is a fact that for max range theta is always equal to 45 degrees...

OpenStudy (anonymous):

sin 2 theta = sin 90 2 theta = 90 theta = 45

OpenStudy (anonymous):

can u solve for it ?

OpenStudy (anonymous):

sin 2 theta = sin 90 how???

OpenStudy (anonymous):

sin 90 = 1. so in place of 1, put sin 90

OpenStudy (anonymous):

sin90=0

OpenStudy (anonymous):

sin 90 = 1 not zero

OpenStudy (anonymous):

oh sorry .....

OpenStudy (radar):

All you have to do is plug the values that you already have to solve. Here is the equation\[H=u ^{2}\sin ^{2}\theta /2g\] \[H=160 ^{2}\sin ^{2}45^{o}/2*10\] H=(25600)(.5)/20 H=12800/20=640 m